OpenStudy (anonymous):

literal questions Solve ab - c = 10 for c lol i dont understand. how do i do this?

OpenStudy (anonymous):

a + b = 12 a + c = 13 b + c = 10 not quite sure if this answers your problem.

OpenStudy (anonymous):

no im pretty sure that is wrong. the answer is c = -10 + ab or c = ab - 10

OpenStudy (anonymous):

Is this in the context of a triangle?

OpenStudy (anonymous):

Just ad ab on both sides...

OpenStudy (anonymous):

add*

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Or you can just add +c on both sides and -10 from both sides.

OpenStudy (anonymous):

oh lol its subtact by ab and multiply by -1 nvm. This questions is about literal equations not values of variables. - sis (the owner of this account)